A boat took 3 hours to travel 24 km with the current. On the return trip, the boat took 5 hours to
travel 24 km against the current. Determine the speed of the current

Answer by ankor@dixie-net.com(22740)

A boat took 3 hours to travel 24 km with the current.
On the return trip, the boat took 5 hours to travel 24 km against the current.
Determine the speed of the current
:
Let c = rate of the current
Let s = boat speed in still water
then
(s+c) = effective speed with the current
and
(s-c) = effective speed against the current
:
Write a distance equation for each way. dist = time * speed
3(s+c) = 24
and
5(s-c) = 24
5s - 5c = 24
:
Simplify the 1st equation, divide both sides by 3 and you have:
s + c = 8
s = (8-c)
Substitute (8-c) for s in the 2nd equation
5(8-c) - 5c = 24
40 - 5c - 5c = 24
40 - 24 = 10c
16 = 10c
c = 16/10
c = 1.6 km/hr is the rate of the current
:
:
You can confirm this for yourself, find s and check in both equations